# FedX: Explanation-Guided Pruning for Communication-Efficient Federated Learning in Remote Sensing

### v1

@@ Abstract
Federated learning (FL) enables the collaborative training of deep neural networks across decentralized data archives (i.e., clients), where each client stores data locally and only shares model updates with a central server. This makes FL a suitable learning paradigm for remote sensing (RS) image classification tasks, where data centralization may be restricted due to legal and privacy constraints. However, a key challenge in applying FL to RS tasks is the communication overhead caused by the frequent exchange of large model updates between clients and the central server. To address this issue, in this paper we propose a novel strategy (denoted as FedX) that uses explanation-guided pruning to reduce communication overhead by minimizing the size of the transmitted models without compromising performance. FedX leverages backpropagation-based explanation methods to estimate the task-specific importance of model components and prunes the least relevant ones at the central server. The resulting sparse global model is then sent to clients, substantially reducing communication overhead. We evaluate FedX on multi-label scene classification using the BigEarthNet-S2 dataset and single-label scene classification using the EuroSAT dataset. Experimental results show the success of FedX in significantly reducing the number of shared model parameters while enhancing the generalization capability of the global model, compared to both unpruned model and state-of-the-art pruning methods. The code of FedX will be available at https://git.tu-berlin.de/rsim/FedX.
Index Terms:
Federated learning, model pruning, image classification, explanation methods, remote sensing.
@@ I Introduction
Remote sensing (RS) image classification plays a crucial role in a wide range of RS applications [1, 2, 3]. In recent years, deep learning (DL) techniques have significantly advanced the performance of RS image classification by enabling the extraction of discriminative features to represent the complex content of images [4, 5]. Traditional DL techniques typically rely on centralized data aggregation. However, RS images are often distributed across decentralized image archives (i.e., clients) [6]. Due to bandwidth and storage limitations, consolidating all RS data on a central server is challenging [7]. This challenge is further compounded due to the continuous increase in the volume and resolution of RS data. In addition to data storage constraints, there are also growing concerns about data sensitivity and privacy in decentralized image archives. In RS, especially with the continued advancement of high-resolution imagery, there is an increasing risk of capturing sensitive information [8]. As an example, satellite images used in urban monitoring can reveal detailed views of private properties, such as building layouts and vehicles, thereby raising ethical and legal concerns related to personal privacy. Additionally, while RS images alone may not raise privacy concerns, their combination with other data sources can lead to situations that pose significant privacy risks. As an example, the fusion of RS images and socio-economic data (e.g., income levels, education, health statistics etc.) can significantly increase the risk of re-identification. This enables highly detailed profiling of specific communities or even individuals. In addition, legal regulations often prohibit the sharing of RS images that involve the monitoring of sensitive sites such as borders or critical infrastructure [9]. These challenges highlight the need for a learning approach that enables collaborative model training across distributed RS image archives, where RS images remain unshared due to privacy concerns and legal constraints.
@@
Federated learning (FL) addresses this need by introducing a decentralized training strategy in which clients collaboratively train models by sharing only model updates with a central server. This enables the joint model training without compromising data confidentiality. Moreover, by keeping data localized, FL alleviates the need for centralized data storage and ensures compliance with legal data handling requirements. However, one notable limitation of FL is the communication overhead associated with frequent exchanges of model updates between clients and the central server. In each training round, model parameters are required to be transmitted back and forth, which can be challenging in environments with limited bandwidth or a large number of participating clients. This is particularly relevant in RS, where clients may include edge devices such as unmanned aerial vehicles or satellites with constrained communication capabilities. This overhead can slow down training and introduce latency, making it challenging to deploy FL in real-time or resource-constrained environments. While the communication overhead in FL has been extensively studied in the computer vision (CV) community (see Section II for a detailed review), relatively few works have addressed this limitation in the context of RS [10, 11, 12]. In [10], a feature-based FL framework is proposed for RS image classification, where clients transmit low-dimensional embeddings or pseudo-weights instead of full models. To this end, image representations are extracted from RS images and shared with the server to enable collaborative training with reduced communication. In [11], a satellite-ground FL framework is introduced to address bandwidth limitations in spaceborne RS. Accordingly, model updates are compressed through progressive weight quantization before being transmitted from satellites to ground stations. In [12], a pruning-based FL method is proposed for RS using unmanned aerial vehicles, where clients apply one-shot filter pruning to remove less important convolutional filters prior to communication. To mitigate communication overhead in FL, existing methods in CV are commonly classified into three main categories: 1) model compression [13, 14, 15, 16]; 2) knowledge distillation [17, 18, 19, 20]; and 3) pruning [21, 22, 23, 24]. Model compression and knowledge distillation are widely used strategies for reducing communication overhead in FL. Model compression-based methods address communication overhead by quantizing the size of model updates before transmission. However, these techniques may lead to a drop in model performance due to the removal or approximation of important parameters [25]. They can also introduce computational overhead, as clients and servers often need to perform additional encoding and decoding steps during training. Knowledge distillation-based methods reduce communication by transmitting soft predictions instead of full model parameters, allowing the server to aggregate knowledge from clients without sharing full models. However, these approaches rely on student models that may fail to match the performance of the teacher models, especially for complex tasks. Moreover, knowledge distillation may still leak sensitive information through the shared logits [26]. Pruning strategies offer advantages over both knowledge distillation and model compression by directly reducing model size without requiring separate student models or extra encoding steps. As a result, pruning maintains model structure and minimizes additional computational overhead.
@@
In this paper, we focus on pruning-based methods and investigate the effectiveness of explanation-guided pruning in the context of FL for RS. We propose an adaptation of backpropagation-based explanation techniques to effectively address the communication overhead in RS applications. Our strategy leverages the explanation signals computed at the central server to assess the relevance of model parameters. Based on the relevance scores, the server performs pruning by removing less important neurons or filters from the global model. The pruned global model is then sent back to the clients for the next round of local training. By performing pruning methods at the server side, our strategy reduces the size of the global model transmitted in each round while preserving task-relevant information. This results in significantly lower communication costs, making the strategy particularly well-suited for large-scale, bandwidth-limited, and privacy-sensitive RS environments. It is worth noting that a preliminary version of our strategy was briefly introduced in [27] with limited experimental evaluation. In this paper, we significantly extend that work by presenting a comprehensive description of the proposed strategy and conducting extensive experiments in the context of multi-label scene classification using the BigEarthNet-S2 dataset and single-label scene classification using the EuroSAT dataset. In addition, we extend the proposed methodology by replacing global pruning with a layer-wise pruning configuration. This change addresses the limitation of global pruning, which disproportionately removes components from deeper layers due to lower average relevance scores caused by relevance conservation. Layer-wise pruning configuration addresses this issue by applying thresholds within each layer, resulting in more balanced pruning and improved model robustness under high sparsity. To the best of our knowledge, it is the first FL strategy that integrates explanation-guided pruning to reduce communication overhead while preserving model performance in image classification tasks.
@@ II-C Pruning-Based FL Algorithms
Pruning-based FL algorithms aim to reduce model complexity by identifying and removing less important parameters such as individual weights, neurons, or convolutional filters. In [21], a pruning-based FL framework (denoted as PruneFL) is introduced to reduce communication and computation costs by adaptively pruning local models during training. To this end, PruneFL estimates the importance of model weights based on their magnitude and prunes less significant parameters on the clients before transmitting updates to the server. In [28], a federated dynamic sparse training framework (denoted as FedDST) maintains a fixed global sparsity throughout training. Each client applies layer-wise magnitude-based pruning at every communication round and restores pruned weights using random initialization or gradient-based criteria. In [22], a layer-wise pruning algorithm (denoted as FedLP) is proposed to address the communication overhead in FL by systematically pruning entire layers or layer components. Accordingly, FedLP employs a unified optimization framework that jointly considers pruning ratio and layer selection, allowing each client to independently adapt the pruned model based on local resource constraints. In [23], a distributed pruning framework (denoted as FedTiny) is introduced for FL that generates highly compressed models suitable for memory- and computation-constrained devices. To this end, FedTiny integrates two core components: an adaptive batch normalization selection module to reduce pruning bias caused by heterogeneous local data distributions, and a lightweight progressive pruning module that incrementally determines layer-wise pruning policies under strict resource constraints. In [29], a dynamic pruning framework (denoted as FedDIP) combines extreme pruning with incremental regularization. During local training, each client performs dynamic pruning by removing low-magnitude weights and then progressively applies an $L_{2}$ regularization to stabilize the pruned structure. In [24], a personalized FL algorithm with adaptive pruning is proposed by introducing a two-stage approach in which the initialization phase rapidly identifies sparse subnetworks, and the adaptive pruning phase customizes models to local data. This design enables efficient training while reducing both communication and computational overhead on edge devices. Existing pruning methods rely on simple heuristics such as weight magnitude and offer limited insight into which parameters are removed. This can result in the loss of task-relevant information. To address this, we propose an explanation-guided pruning strategy that uses relevance scores to retain the most informative components during pruning.
@@ III-B Explanation-Guided Pruning at the Server
The main idea of the proposed pruning strategy is to systematically identify and remove globally redundant or less important components of the model based on their quantified task relevance. The global model $\phi^{*}$ is structurally decomposed into a set of $p$ discrete components $\{\psi_{1},\psi_{2},...,\psi_{p}\}$. The granularity of these components is flexible and can be adapted to suit the model architecture and pruning objectives. Specifically, components may correspond to convolutional filters, neurons, or attention heads, each representing a meaningful functional block within the model. This modular representation allows for pruning at various levels of abstraction, from coarse-grained units (e.g., entire layers) to fine-grained elements (e.g., individual filters). To assess the relevance of each component, the server utilizes a reference set denoted as $D_{\text{ref}}=\{(\boldsymbol{x}^{z},\boldsymbol{y}^{z})\}_{z=1}^{M_{ref}}$, where $M_{ref}$ is the number of samples in the reference set. This dataset provides the necessary context for evaluating the contribution of individual components. To ensure that the relevance assessment remains both privacy-preserving and generalizable, $D_{\text{ref}}$ can be constructed using publicly available or proxy data sources. For each input sample $I_{z}\in D_{\text{ref}}$, the explanation-guided pruning methods are applied to propagate the prediction of the model backward through the network to compute relevance scores at the component level. Specifically, for component $k$ in layer $l$, the relevance score $R_{k}^{(l)}(I_{z})$ is determined by the proportional redistribution of relevance from the subsequent layer:
@@
where $v_{k,j}$ reflects the contribution of component $k$ to component $j$ in the next layer, and $R_{j}^{(l+1)}(I_{z})$ is the relevance score at the higher layer. Following component-level relevance estimation, the next step aggregates relevance scores at the component level. This process is repeated for all samples in $D_{\text{ref}}$. Then, the average relevance score of component $\psi_{c}$ across $D_{\text{ref}}$ is computed to obtain its global relevance estimate as follows:
@@
This ensures that the least relevant fraction $q$ of components in each layer is pruned. The choice of $q$ directly affects the trade-off between communication efficiency and generalization capability. A higher $q$ results in stronger pruning, reducing communication overhead while potentially degrading performance. Conversely, a lower $q$ preserves more parameters and maintains higher model performance, at the cost of reduced communication efficiency. After computing the pruning mask, we apply it to the global model parameters:
@@ IV-B Design of Experiments
In the experiments, as a FL algorithm, we exploited FedAvg [35], which updates the global model by aggregating the weighted averages of locally trained models on clients. This algorithm was selected as the aggregation method, as it serves as a widely adopted baseline, allowing fair and interpretable comparisons between different pruning strategies. This choice allows us to ensure that performance differences are caused by the pruning method instead of the aggregation process. In the experiments, the number of communication rounds was set to 20, with each client performing 3 local training epochs per round to allow sufficient model updates while maintaining a manageable overall training duration. During the experiments, we observed that increasing the number of communication rounds or local epochs beyond this point did not lead to notable changes in model performance. We tested our strategy in the context of two learning tasks: 1) multi-label scene classification using BigEarthNet-S2; and 2) single-label scene classification using EuroSAT. We considered six different DL model architectures: ResNet6 [36] ($\approx 1.7$M parameters), ResNet18 [36] ($\approx 11.7$M parameters), ResNet50 [36] ($\approx 25.5$M parameters), ViT-B16 [37] ($\approx 86$M parameters), ConvNeXt V2 Atto [38] ($\approx 3.7$M parameters), and ConvNeXt V2 Nano [38] ($\approx 14.0$M parameters). As explanation-guided pruning methods, we selected LRP [39] and IG [40] due to their proven success. We extended their applicability to FL by computing relevance scores to identify less important parameters and applying structured pruning to the global model before redistributing it to clients. For LRP hyperparameterization, each model was divided into four consecutive parts following the approach in [40]. For the ResNet architectures, we used the epsilon rule in the low-level hidden layers, the gamma rule in the mid-level and high-level hidden layers, and the epsilon rule again in the fully connected layers. For ViT-B16, we employed the epsilon rule uniformly in all four parts. We considered layer-wise pruning configuration, where model parameters are pruned independently within each layer, and a global pruning configuration, where the entire model is pruned at once. All models were trained using the Adam optimizer with a learning rate of 0.001, weight decay of 1e-4, and a batch size of 512. Pruning was applied after a warmup phase of 10 communication rounds ($\upsilon=10$). This choice reflects the trade-off between achieving model stability and minimizing communication overhead. Performance was evaluated using micro mean Average Precision (mAP) for the multi-label classification task and overall accuracy for the single-label classification task. We used mAP as it is well suited for multi-label classification tasks with class imbalance, providing a comprehensive measure by aggregating true positives, false positives, and false negatives across all classes.

### v2

@@ Abstract
Federated learning (FL) enables the collaborative training of deep neural networks across decentralized data archives (i.e., clients), where each client stores data locally and only shares model updates with a central server. This makes FL a suitable learning paradigm for remote sensing (RS) image classification tasks, where data centralization may be restricted due to legal and privacy constraints. However, a key challenge in applying FL to RS tasks is the communication overhead caused by the frequent exchange of large model updates between clients and the central server. To address this issue, in this paper we propose a novel strategy (denoted as FedX) that uses explanation-guided pruning to reduce communication overhead by minimizing the size of the transmitted models without compromising performance. FedX leverages backpropagation-based explanation methods to estimate the task-specific importance of model components and prunes the least relevant ones at the central server. The resulting sparse global model is then sent to clients, substantially reducing communication overhead. We evaluate FedX on multi-label scene classification using the BigEarthNet-S2 dataset and single-label scene classification using the EuroSAT dataset. Experimental results show the success of FedX in significantly reducing the number of shared model parameters while enhancing the generalization capability of the global model, compared to both unpruned model and state-of-the-art pruning methods. The code of FedX is publicly available at https://git.tu-berlin.de/rsim/FedX.
@@ I Introduction
Remote sensing (RS) image classification plays a crucial role in a wide range of RS applications [25, 20, 43]. In recent years, deep learning (DL) techniques have significantly advanced the performance of RS image classification by enabling the extraction of discriminative features to represent the complex content of images [33, 10]. Most existing DL techniques typically require that all training data to be collected and stored on a single central server for model training. However, RS images are often distributed across decentralized image archives (i.e., clients) [4]. Due to bandwidth and storage limitations, consolidating all RS data on a central server is challenging [5]. This challenge is further compounded due to the continuous increase in the volume and resolution of RS data. In addition to data storage constraints, there are also growing concerns about data sensitivity and privacy in decentralized image archives. In RS, especially with the continued advancement of high-resolution imagery, there is an increasing risk of capturing sensitive information [28]. As an example, satellite images used in urban monitoring can reveal detailed views of private properties, such as building layouts and vehicles, thereby raising ethical and legal concerns related to personal privacy. Additionally, while RS images alone may not raise privacy concerns, their combination with other data sources can lead to situations that pose significant privacy risks. As an example, the fusion of RS images and socio-economic data (e.g., income levels, education, health statistics etc.) can significantly increase the risk of re-identification. This enables highly detailed profiling of specific communities or even individuals. In addition, legal regulations often prohibit the sharing of RS images that involve the monitoring of sensitive sites such as borders or critical infrastructure [1]. These challenges highlight the need for a learning approach that enables collaborative model training across distributed RS image archives, where RS images remain unshared due to privacy concerns and legal constraints.
@@
Federated learning (FL) addresses this need by introducing a decentralized training strategy in which clients collaboratively train models by sharing only model updates with a central server. This enables the joint model training without compromising data confidentiality. Moreover, by keeping data localized, FL alleviates the need for centralized data storage and ensures compliance with legal data handling requirements. However, one notable limitation of FL is the communication overhead associated with frequent exchanges of model updates between clients and the central server. In each training round, model parameters are required to be transmitted back and forth, which can be challenging in environments with limited bandwidth or a large number of participating clients. This is particularly relevant in RS, where clients may include edge devices such as unmanned aerial vehicles or satellites with constrained communication capabilities. This overhead can slow down training and introduce latency, making it challenging to deploy FL in real-time or resource-constrained environments. While the communication overhead in FL has been extensively studied in the computer vision (CV) community (see Section II for a detailed review), relatively few works have addressed this limitation in the context of RS [9, 39, 31, 22, 15, 42].
In [9], a feature-based FL framework is proposed for RS image classification, where clients transmit low-dimensional embeddings or pseudo-weights instead of full models. To this end, image representations are extracted from RS images and shared with the server to enable collaborative training with reduced communication. In [39], a satellite-ground FL framework is introduced to address bandwidth limitations in spaceborne RS. Accordingly, model updates are compressed through progressive weight quantization before being transmitted from satellites to ground stations. In [31], a pruning-based FL method is proposed for RS using unmanned aerial vehicles, where clients apply one-shot filter pruning to remove less important convolutional filters prior to communication. In [22], a diffusion model–driven FL framework (denoted as FedDiff) is proposed for secure multi-modal RS data fusion across multiple clients. To this end, a dual-branch diffusion architecture with a lightweight communication module is embedded into the FL framework to enable communication-efficient and privacy-preserving multi-modal knowledge exchange. In [15], a federated generative adversarial framework for privacy-preserving synthesis of multi-source RS time series (denoted as FedScoreGAN) is proposed under non-IID data. The framework integrates score-based generative modeling with federated aggregation and incorporates a communication-aware design via parameter sparsification. In [42], a federated DL scheme with prototype matching (denoted as FedPM) is proposed to enable training across multiple clients with heterogeneous data. The scheme enforces class-wise prototype alignment during local optimization to reduce feature distribution discrepancies across clients. Prototype deviations are then leveraged to guide attention-weighted global aggregation, while sparse ternary compression is integrated to alleviate communication overhead. To mitigate communication overhead in FL, existing methods in CV are commonly classified into three main categories: 1) model compression [30, 23, 44, 35]; 2) knowledge distillation [26, 11, 41, 37]; and 3) pruning [17, 46, 16, 45]. Model compression and knowledge distillation are widely used strategies for reducing communication overhead in FL. Model compression-based methods address communication overhead by quantizing the size of model updates before transmission. However, these techniques may lead to a drop in model performance due to the removal or approximation of important parameters [7]. They can also introduce computational overhead, as clients and servers often need to perform additional encoding and decoding steps during training. Knowledge distillation-based methods reduce communication by transmitting soft predictions instead of full model parameters, allowing the server to aggregate knowledge from clients without sharing full models. However, these approaches rely on student models that may fail to match the performance of the teacher models, especially for complex tasks. Moreover, knowledge distillation may still leak sensitive information through the shared logits [38]. Pruning strategies offer advantages over both knowledge distillation and model compression by directly reducing model size without requiring separate student models or extra encoding steps. As a result, pruning maintains model structure and minimizes additional computational overhead.
@@
In this paper, we focus on pruning-based methods and investigate the effectiveness of explanation-guided pruning in the context of FL for RS. To address this, we propose a strategy that utilizes backpropagation-based explanation methods to effectively reduce the communication overhead in FL for RS applications. We specifically focus on explanation-guided pruning in FL because magnitude-based and locally driven pruning methods can be less reliable for identifying globally relevant components. After aggregation across heterogeneous clients, updates from different data distributions may partially cancel each other out, causing parameters that are important for some clients to have small magnitudes in the global model despite carrying task relevant information. To overcome this limitation, our strategy leverages the explanation signals computed at the central server to assess the relevance of model parameters. Based on the relevance scores, the server performs pruning by removing less important neurons or filters from the global model. The pruned global model is then sent back to the clients for the next round of local training. By performing pruning methods at the server side, our strategy reduces the size of the global model transmitted in each round while preserving task-relevant information. This results in significantly lower communication costs, making the strategy particularly well-suited for large-scale, bandwidth-limited, and privacy-sensitive RS environments. It is worth noting that a preliminary version of our strategy was briefly introduced in [19] with limited experimental evaluation. In this paper, we significantly extend that work by presenting a comprehensive description of the proposed strategy and conducting extensive experiments in the context of multi-label scene classification using the BigEarthNet-S2 dataset and single-label scene classification using the EuroSAT dataset. In addition, we extend the proposed methodology by replacing global pruning with a layer-wise pruning configuration. This change addresses the limitation of global pruning, which disproportionately removes components from deeper layers due to lower average relevance scores caused by relevance conservation. Layer-wise pruning configuration addresses this issue by applying thresholds within each layer, resulting in more balanced pruning and improved model robustness under high sparsity. To the best of our knowledge, it is the first FL strategy that utilizes explanation-guided pruning to reduce communication overhead while preserving model performance in image classification tasks.
@@ II-C Pruning-Based FL Algorithms
Pruning-based FL algorithms aim to reduce model complexity by identifying and removing less important parameters such as individual weights, neurons, or convolutional filters. In [21], a pruning-based FL framework (denoted as LotteryFL) utilizes the Lottery Ticket Hypothesis to identify sparse subnetworks on each client. Each client independently learns a lottery ticket network, and only these client-specific subnetworks are communicated between the server and the participating clients during training. In [17], a pruning-based FL framework (denoted as PruneFL) is introduced to reduce communication and computation costs by adaptively pruning local models during training. To this end, PruneFL estimates the importance of model weights based on their magnitude and prunes less significant parameters on the clients before transmitting updates to the server. In [3], a federated dynamic sparse training framework (denoted as FedDST) maintains a fixed global sparsity throughout training. Each client applies layer-wise magnitude-based pruning at every communication round and restores pruned weights using random initialization or gradient-based criteria. In [46], a layer-wise pruning algorithm (denoted as FedLP) is proposed to address the communication overhead in FL by systematically pruning entire layers or layer components. Accordingly, FedLP employs a unified optimization framework that jointly considers pruning ratio and layer selection, allowing each client to independently adapt the pruned model based on local resource constraints. In [16], a distributed pruning framework (denoted as FedTiny) is introduced for FL that generates highly compressed models suitable for memory- and computation-constrained devices. To this end, FedTiny integrates two core components: an adaptive batch normalization selection module to reduce pruning bias caused by heterogeneous local data distributions, and a lightweight progressive pruning module that incrementally determines layer-wise pruning policies under strict resource constraints. In [27], a dynamic pruning framework (denoted as FedDIP) combines extreme pruning with incremental regularization. During local training, each client performs dynamic pruning by removing low-magnitude weights and then progressively applies an $L_{2}$ regularization to stabilize the pruned structure. In [45], a personalized FL algorithm with adaptive pruning is proposed by introducing a two-stage approach in which the initialization phase rapidly identifies sparse subnetworks, and the adaptive pruning phase customizes models to local data. This design enables efficient training while reducing both communication and computational overhead on edge devices. Existing pruning methods rely on simple heuristics such as weight magnitude and offer limited insight into which parameters are removed. This can result in the loss of task-relevant information. To address this, we propose an explanation-guided pruning strategy that uses relevance scores to retain the most informative components during pruning.
@@ III-B Explanation-Guided Pruning at the Server
The main idea of the proposed pruning strategy is to systematically identify and remove globally redundant or less important components of the model based on their quantified task relevance. The global model $\phi^{*}$ is structurally decomposed into a set of $p$ discrete components $\{\psi_{1},\psi_{2},...,\psi_{p}\}$. Each component $\psi_{c}$ corresponds to a structured unit of the network (e.g., a convolutional filter), enabling structured pruning and communication-efficient model updates. The granularity of these components is flexible and can be adapted to suit the model architecture and pruning objectives. Specifically, components may correspond to convolutional filters, neurons, or attention heads, each representing a meaningful functional block within the model. This modular representation allows for pruning at various levels of abstraction, from coarse-grained units (e.g., entire layers) to fine-grained elements (e.g., individual filters). To assess the relevance of each component, the server utilizes a reference set denoted as $D_{\text{ref}}=\{(\boldsymbol{x}^{z},\boldsymbol{y}^{z})\}_{z=1}^{M_{ref}}$, where $M_{ref}$ is the number of samples in the reference set. This dataset provides the necessary context for evaluating the contribution of individual components. To ensure that the relevance assessment remains both privacy-preserving and generalizable, $D_{\text{ref}}$ can be constructed using publicly available or proxy data sources. For each input sample $I_{z}\in D_{\text{ref}}$, the explanation-guided pruning methods are applied to propagate the prediction of the model backward through the network to compute relevance scores at the component level. Specifically, for component $k$ in layer $l$, the relevance score $R_{k}^{(l)}(I_{z})$ is determined by the proportional redistribution of relevance from the subsequent layer:
@@
where $v_{k,j}$ reflects the contribution of component $k$ to component $j$ in the next layer, and $R_{j}^{(l+1)}(I_{z})$ is the relevance score at the higher layer. The relevance is computed in a layer-wise backpropagation manner from the output to the input layer using specific propagation rules that ensure relevance conservation throughout the network [29]. For the final layer $L$, the relevance $R_{k}^{(L)}(I_{z})$ is initialized as the output score corresponding to the target class $c$. Following component-level relevance estimation, the next step aggregates relevance scores at the component level. This process is repeated for all samples in $D_{\text{ref}}$. Then, the average relevance score of component $\psi_{c}$ across $D_{\text{ref}}$ is computed to obtain its global relevance estimate as follows:
@@
The binary mask $\mathcal{M}(\psi_{c})\in\{0,1\}$ indicates whether component $\psi_{c}$ is retained ($\mathcal{M}(\psi_{c})=1$) or removed ($\mathcal{M}(\psi_{c})=0$). This ensures that the least relevant fraction $q$ of components in each layer is pruned. The choice of $q$ directly affects the trade-off between communication efficiency and generalization capability. A higher $q$ results in stronger pruning, reducing communication overhead while potentially degrading performance. Conversely, a lower $q$ preserves more parameters and maintains higher model performance, at the cost of reduced communication efficiency. After computing the pruning mask, we apply it to the global model parameters:
@@ IV-B Design of Experiments
In the experiments, as a FL algorithm, we exploited FedAvg [24], which updates the global model by aggregating the weighted averages of locally trained models on clients. This algorithm was selected as the aggregation method, as it serves as a widely adopted baseline, allowing fair and interpretable comparisons between different pruning strategies. This choice allows us to ensure that performance differences are caused by the pruning method instead of the aggregation process. In the experiments, the number of communication rounds was set to 20, with each client performing 3 local training epochs per round to allow sufficient model updates while maintaining a manageable overall training duration. During the experiments, we observed that increasing the number of communication rounds or local epochs beyond this point did not lead to notable changes in model performance. We tested our strategy in the context of two learning tasks: 1) multi-label scene classification using BigEarthNet-S2; and 2) single-label scene classification using EuroSAT. We considered six different DL model architectures: ResNet6 [13] ($\approx 1.7$M parameters), ResNet18 [13] ($\approx 11.7$M parameters), ResNet50 [13] ($\approx 25.5$M parameters), ViT-B16 [8] ($\approx 86$M parameters), ConvNeXt V2 Atto [36] ($\approx 3.7$M parameters), and ConvNeXt V2 Nano [36] ($\approx 14.0$M parameters). As explanation-guided pruning methods, we selected LRP [40] and IG [12] due to their proven success and reliability in RS scene classification [18]. We extended their applicability to FL by computing relevance scores to identify less important parameters and applying structured pruning to the global model before redistributing it to clients. For LRP hyperparameterization, each model was divided into four consecutive parts following the approach in [12]. For the ResNet architectures, we used the epsilon rule in the low-level hidden layers, the gamma rule in the mid-level and high-level hidden layers, and the epsilon rule again in the fully connected layers. For ViT-B16, we employed the epsilon rule uniformly in all four parts. The relevance scores are computed only at the central server using a publicly available reference dataset $D_{ref}$, which is independent of the local data on clients. In our experiments, $D_{ref}$ is constructed by randomly sampling 500 images from the validation set. This reference set is used exclusively for server-side relevance estimation and pruning-mask generation, and it is not used for client-side training. Since the relevance computation is performed entirely at the central server, local data is not shared with the central server. We considered layer-wise pruning configuration, where model parameters are pruned independently within each layer, and a global pruning configuration, where the entire model is pruned at once. All models were trained using the Adam optimizer with a learning rate of 0.001, weight decay of 1e-4, and a batch size of 512. Pruning was applied after a warmup phase of 10 communication rounds ($\upsilon=10$). This choice reflects the trade-off between achieving model stability and minimizing communication overhead. Performance was evaluated using micro mean Average Precision (mAP) for the multi-label classification task and overall accuracy for the single-label classification task. We used mAP as it is well suited for multi-label classification tasks with class imbalance, providing a comprehensive measure by aggregating true positives, false positives, and false negatives across all classes.